# Graph-based RAG Enhancement via Global Query Disambiguation and Dependency-Aware Reranking

### v1

@@ 4.1 Experimental Setup
Table 2: The QA performance across two datasets and four key metrics: answer faithfulness, context recall, and context precision.
@@
GraphRAG 69.56% 86.80% 62.06% 61.18% 68.94% 69.73% 61.57% 60.40% 69.25% 0.78.27% 61.82% 60.70%
LightRAG 72.40% 86.78% 65.18% 67.00% 67.92% 68.68% 60.99% 58.42% 68.66% 77.73% 63.08% 0.6271%
PankRAG 0.7479 0.9653 0.8789 0.8400 0.7062 0.8875 0.6499 0.6535 0.7271 0.9264 0.7644 0.7468
4.1.1 Datasets
We adopt the experimental framework of recent graph-based RAG studies and evaluate our method on six real-world datasets, encompassing both SCQ and ACQ. We use the Musique and MultiHop-RAG datasets for SCQ, which require reasoning across multiple passages, understanding implicit relationships, and synthesizing information. ACQ demands holistic curation, multi-perspective synthesis, and coherent organization of information at scale. We evaluate our method on the UltraDomin dataset Qin et al. (2024), which is drawn from 428 college-level textbooks across 18 domains, selecting the four domains used (Agriculture, Legal, Computer Science, and Mix), whose detailed information is as follows in LightRAG’s evaluation. We follow the standardized preprocessing pipelines from GraphRAG and LightRAG.
Figure 3: Impact of a pair of hyperparameters in PankRAG on the Mix(left) and Agriculture(right) dataset
4.1.2 Baselines
We compare PathRAG with the following state-of-the-art methods across all datasets:
  • NaiveRAG Gao et al. (2024): In this RAG method, text is divided into chunks, stored as embeddings in a vector database, and the most semantically similar chunks are retrieved during queries to enable efficient matching.
  • GraphRAG Edge et al. (2024): In this graph-based RAG method, they hierarchically organize knowledge through entity-level descriptions and clustered community summaries. For global queries, it retrieves community-focused knowledge and aggregates related entities to produce comprehensive responses.
  • LightRAG Guo et al. (2024): In this graph-based RAG method, graph structures are integrated with vector representations, and a dual-level retrieval paradigm is employed to enhance both the efficiency and relevance of retrieval.
4.1.3 Evaluation Metrics
  • SCQ Dataset In the SCQ dataset, we evaluate all models using the Ragas Es et al. (2024) framework across four critical metrics: answer relevance, faithfulness, context recall, and context precision.
  • ACQ Dataset In the absence of ground-truth answers, we adopt LLM-based evaluation protocols following prior work. Specifically, we leverage GPT-4o-mini to assess generated answers across six dimensions: Comprehensiveness, Diversity and Empowerment, adopted from LightRAG Guo et al. (2024), and three novel dimensions—Logicality, Relevance, and Coherence-introduced by recent LLM evaluation research Chen et al. (2025). Answers generated by all baselines and our method are pairwise compared using win-rate analysis, where a higher win rate reflects superior performance relative to alternatives. To mitigate presentation bias, answer order is systematically alternated, with final results reporting averaged win rates. How much detail does the answer provide to cover all aspects and details of the question?
@@ 4.2 Overall Performance on the ACQ Datasets
As demonstrated in Table 1, PathRAG consistently surpasses baseline methods across all evaluation metrics and datasets.
From the perspective of benchmark datasets, PankRAG achieves average win rates of 72.47% on the Agriculture dataset, 69.87% on the Legal dataset, 67.77% on the CS dataset, and 68.4% on the Mix dataset. These results demonstrate PankRAG’s generalizability in effectively facilitating LLMs to capture hidden critical information and relationships in ACQ across datasets of varying scales, thereby achieving robust and superior performance.
From the evaluation dimension perspective, PankRAG attains average win rates of 65.55% in comprehensiveness, 69.64% in diversity, 72.33% in logicality, 68.5% in relevance, 73.00% in coherence, and 68.86% in empowerment. This consistent outperformance across all dimensions, attributed to the globally-aware pathway resolution and dependency-aware reranking mechanisms, underscores PankRAG’s comprehensive advantages over baseline RAG methods.
Figure 4: Impact of a pair of hyperparameters in PankRAG on the Multihop-RAG dataset
@@ 4.3 Overall Performance on the SCQ Datasets
As demonstrated in LABEL:tab:specificq_res, PankRAG consistently achieves the best performance across all experimental settings.
Compared to NaiveRAG, GraphRAG, and LightRAG, PankRAG achieves average metric improvements of 17.42%, 11.61%, 11.08%. Averaged over multiple datasets, the results are as follows. PankRAG achieves approximately 5.99% higher than NaiveRAG, 3.46% higher than GraphRAG, and 4.05% higher than LightRAG in Answer Relevance; 25.94% higher than NaiveRAG, 14.37% higher than GraphRAG, and 14.91% higher than LightRAG in Faithfulness; 19.15% higher than NaiveRAG, 14.62% higher than GraphRAG, and 13.36% higher than LightRAG in Context Recall; and 18.57% higher than NaiveRAG, 13.98% higher than GraphRAG, and 11.97% higher than LightRAG in Context Precision. These results demonstrate PankRAG’s superiority over prior graph-based RAG methods in handling concrete complex problems.
Notably, GraphRAG and LightRAG show minimal improvements in Context Recall and Context Precision (4%-6% over NaiveRAG), indicating negligible enhancement in retrieved context quality. This aligns with the identified challenge that LLMs struggle to interpret hidden relationships or information in SCQ, leading to misaligned retrieval targets. While these methods enhance retrieval via graph structures, their limited context quality gains highlight unresolved limitations. In contrast, PankRAG’s significant improvements in Context Recall/Precision suggest effective mitigation of this issue, attributed to its globally-aware pathway resolution generation module and dependency-aware reranking mechanism. It enables LLMs to better capture hidden relationships and thereby improve context relevance. Concurrent enhancements in Faithfulness and Answer Relevance further indicate reduced hallucination risks and higher answer quality, underscoring PankRAG’s comprehensive advancements over existing graph-based RAG approaches.
Table 3: Ablation study on abstract queries on the global-aware pathway resolution generation module and dependency-aware reranking mechanism in PankRAG for Abstract Queries.
@@
PankRAG - Pank PankRAG - Ank PankRAG - Pank PankRAG - Ank PankRAG - Pank PankRAG - Ank PankRAG - Pank PankRAG - Ank
@@
PankRAG - Ank PankRAG PankRAG - Ank PankRAG PankRAG - Ank PankRAG PankRAG - Ank PankRAG
@@
Table 4: Ablation study on abstract queries on the global-aware pathway resolution generation module and dependency-aware reranking mechanism in PankRAG for Specific Queries.
Multihop-RAG MuSiQue Average
PankRAG-Pank 67.96 88.21% 62.94 62.21 67.45% 75.16% 59.90% 58.73% 67.70% 81.69% 61.42% 60.47%
PankRAG 0.7479 0.9653 0.8789 0.8400 0.7062 0.8875 0.6499 0.6535 0.7271 0.9264 0.7644 0.7468
@@ 4.4 Hyperparameter Analysis
The parameters $\alpha$ and $\beta$ are designed to optimize the original quality score by integrating a novel similarity metric, constrained by $\alpha$ + $\beta$ = 1 to ensure the normalized score range. Experimental results demonstrate that the optimal parameters for SCQ are $\alpha$ > 0.6 $\beta$ < 0.4, whereas for ACQ, $\alpha$ = 0.75 and $\beta$ = 0.25 , with both configurations maximizing retrieval performance.
In SCQ datasets, we evaluated retrieval performance across various ($\alpha$:$\beta$) ratios in Figure 4. As ($\alpha$:$\beta$) increases, key metrics—including answer relevance, context recall, and context precision—monotonically improve, peaking at ($\alpha$:$\beta$) = 0.6:0.4, before declining. Although faithfulness attains its maximum at ($\alpha$:$\beta$) = 0.4:0.6, its value at $0.6:0.4$ is only 0.0027 lower, indicating a negligible trade-off. Thus, $\alpha$ = 0.6 and $\beta$ = 0.4 represent the optimal balance. We indicate that when $\alpha$ > 0.6 $\beta$ < 0.4, the weight on sub-question answers is insufficient to adequately reinforce and validate retrieved context. Conversely, when $\alpha$ < 0.6 $\beta$ > 0.4, over-emphasis on sub-question answers introduces negative hallucinations that outweigh positive reinforcement, degrading performance. At $\alpha$ = 0.6 and $\beta$ = 0.4, these opposing effects are balanced, yielding the best overall retrieval results.
In ACQ datasets, We adjust a pair of hyperparameters on the Mix and Agriculture dataset, and then calculate the win rates compared with LightRAG, the best baseline. As shown in  Figure 3, parameter configuration of $\alpha$ = 0.75 and $\beta$ = 0.25 achieves the highest win rates across the majority of the six dimensions and secures the second-highest performance in all remaining cases. For ACQ, the optimal parameter configuration that maximizes retrieval performance is achieved when $\alpha$ = 0.75 and $\beta$ = 0.25.
@@ 4.5 Ablation Studies
In ACQ datasets, Since our proposed dependency-aware reranking mechanism (Ank) operates on top of the globally-aware pathway resolution (Path), we first evaluated the effectiveness of Path by measuring the win rates of PankRAG-Path versus PankRAG-Ank across six dimensions, as shown in the Table 3. Subsequently, to validate the synergistic enhancement of Ank on the Path-integrated graph RAG framework, we compared the win rates of PankRAG-Ank against PankRAG across the same six dimensions.
As shown in Table 3, PathRAG achieves an average win rate of 57.75% against PathRAG-Ank, PathRAG-Ank attains a 55.27% average win rate over PathRAG-Pank. These results demonstrate the necessity of the globally-aware pathway resolution and further validate the dependency-aware reranking mechanism’s critical role in optimizing RAG systems incorporating globally-aware pathway resolution.
In SCQ datasets, we separately evaluated PankRAG-Pank, PankRAG-Ank, and PankRAG across four metrics to validate the effectiveness of the globally-aware pathway resolution (Path) and the dependency-aware rerank (Ank), respectively.
As shown in the Table 4, PankRAG surpasses PankRAG-Ank by approximately 3.1% and PankRAG-Pank by 4.51% in Answer Relevance. For Faithfulness, PankRAG outperforms PankRAG-Ank by 4.85% and PankRAG-Pank by 10.96%. In Context Precision, PankRAG exceeds PankRAG-Ank by 3.35% and PankRAG-Pank by 15.22%, while in Context Recall, it surpasses PankRAG-Ank by 6.05% and PankRAG-Pank by 14.21%. Notably, PankRAG-Ank achieves significant improvements over PankRAG-Pank in Context Precision and Recall, indicating enhanced retrieval performance and underscoring the necessity of the globally-aware pathway resolution. Concurrently, PankRAG’s consistent superiority over PankRAG-Ank across all metrics highlights the critical role of the dependency-aware reranking mechanism.
@@ 5 Conclusion
In this paper, we propose PankRAG, a novel RAG framework designed to enhance LLMs’ capacity to accurately interpret and retrieve information for complex queries. PankRAG combines a globally aware pathway-resolution module with a dependency-aware reranking mechanism. By executing globally coherent reasoning pathways and augmenting retrieved content through dependency-aware reranking, our framework captures and resolves hidden relationships and information in both abstract and specific complex queries, significantly enhancing retrieval performance and response quality.
Limitations
While PankRAG enhances retrieval accuracy and comprehensiveness through globally-aware query resolution paths and dependency-aware reranking mechanisms, it faces two key limitations. First, resolving multiple sub-problems during retrieval necessitates additional LLM invocations, which improves quality of response but introduces higher computational overhead. Second, our evaluation follows the relative metrics of existing graph-based RAG methods, which may lack stability compared to absolute benchmarks. Future work will focus on establishing more reliable and absolute evaluation metrics tailored to graph-based RAG frameworks.
Appendix A Appendix
A.1 Datasets
A.1.1 ACQ
  • CS: This dataset focuses on core computer science domains, emphasizing data science and software engineering. It highlights practical implementations in machine learning (e.g., recommendation systems, classification algorithms) and big data processing (e.g., Spark-based real-time analytics).
  • Legal: This dataset centers on corporate legal frameworks, covering restructuring procedures, compliance protocols, legal agreements, and governance mechanisms, with applications in legal-financial sector contexts.
  • Agriculture: This dataset addresses agricultural practices, including beekeeping strategies, crop production optimization, disease prevention techniques, and farm management methodologies.
  • Mix: This dataset compiles literary classics across genres (essays, poetry, biographies), spanning philosophy, history, and cultural studies through interdisciplinary perspectives.
A.2 Evaluation
A.2.1 ACQ
LLM evaluate RAG responses according to following seven dimensions:
  • Comprehensiveness: How much detail does the answer provide to cover all aspects and details of the question?
  • Diversity: How varied and rich is the answer in providing different perspectives and insights on the question?
  • Logicality: How logically does the answer respond to all parts of the question?
  • Relevance: How relevant is the answer to the question, staying focused and addressing the intended topic or issue?
  • Coherence: How well does the answer maintain internal logical connections between its parts, ensuring a smooth and consistent structure?
  • Empowerment: How well does the answer help the reader understand and make informed judgments about the topic?
To mitigate ordering effects in comparative evaluations, we implement a systematic rotation of answer positions between the two methods under investigation. Each method’s output is alternately assigned to the primary position (Answer A) and the secondary position (Answer B) across evaluation trials. This dual-perspective counterbalancing enables the computation of bidirectional win rates, ensuring evaluation parity while controlling for positional bias inherent in sequential prompt structures.
A.2.2 SCQ
  • This metric quantifies the degree to which generated responses directly address the original query {r_q}, emphasizing completeness and specificity while penalizing redundant or tangential content. Automated evaluation involves generating {r_n} hypothetical questions from the answer using LLMs, embedding both original and generated questions (via text-embedding-ada-002), and computing their cosine similarity. The final score is formally defined as:
    $\text{AR}=\frac{1}{n}\sum_{i=1}^{n}\cos(\mathbf{e}_{q},\mathbf{e}_{q_{i}})$ (6)
    where {e_q} denotes the original question and {e_q_i} represent LLM-generated counterpart questions.
  • Faithfulness(F): This metric measures the factual grounding of answers within provided contexts, critical for mitigating hallucinations. The evaluation involves: 1. Decomposing answers into atomic claims from 2. Verifying each claim against through entailment checks. The faithfulness score is calculated as:
    $F=\frac{|\text{Verified Claims}|}{|\text{Total Claims}|}$ (7)
    where {_V} denotes verified claims and {_S} the total extracted statements.
  • Context Recall(CR): This metric evaluates the precision of retrieved contexts by quantifying the proportion of context sentences directly relevant to answering the query. Using LLM extraction, critical sentences {S_ext} through structured prompting. The context are identified from {C_q} recall score (CR) is derived as:
    $\text{CR}=\frac{|S_{\text{ext}}|}{|c(q)|}$ (8)
    where {c(q)} represents the total sentences in the context. represents the total sentences in the context. Low CR scores indicate excessive irrelevant content, highlighting retrieval inefficiency.
A.3 Case Study
To demonstrate the advantages of PankRAG over existing RAG systems, we conduct a comparative analysis between PankRAG and baseline methods using the ACQ task ("How can theoretical principles from programming language theory be made accessible to new developers?") from the CS dataset. Due to space constraints, we present only the case study of the ACQ task with PankRAG versus LightRAG. As shown in the table, with the help of a hierarchical resolution path with global awareness in Figure 5, PankRAG outperforms LightRAG across all six evaluation metrics, with detailed analyses evaluated from LLMs provided below.
LightRAG Response Several approaches can facilitate the understanding of these principles, making them more relatable and useful for beginners embarking on their coding journeys. 1. Educational Resources : […] 2. Hands-On Experience : […] 3. Use of Simplified Language and Analogies : […] 4. Community Building and Support : […] Conclusion: Integrating educational resources, hands-on experience, simplified explanations, supportive communities, new developers can better access the theoretical principles of programming language theory…
PankRAG Response Making Theoretical Principles Accessible to New Developers To facilitate a smoother transition for new developers into the complexities of programming language theory … A structured approach that incorporates simplifications, practical engagement, interactive learning resources, and collaborative support will enhance comprehension and retention. 1. Simplification of Concepts : […] 2. Practical Engagement : […] 3. Interactive Learning Resources : […] 4. Collaborative Learning and Support : […] Conclusion: In summary, to make theoretical principles from programming language theory accessible to new developers, a multifaceted approach that emphasizes simplification, practical engagement, interactive resources, and collaborative support…
LLM Evaluation:
  • Comprehensiveness: “Winner”: "Answer2"
    “Explanation”: Answer2(PankRAG) provides a detailed breakdown of strategies to make theoretical principles accessible, including simplification, practical engagement, interactive resources, and collaborative support. It covers various aspects of the topic thoroughly, while Answer1(LightRAG) …
  • Diversity: “Winner”: "Answer2"
    “Explanation”: Answer2(PankRAG) presents a wider range of strategies and methods, including mentorship programs and specific examples of interactive tools. It emphasizes diverse learning approaches, while Answer1(LightRAG) …
  • Logicality: “Winner”: "Answer2"
    “Explanation”: Answer2(PankRAG) follows a clear, structured approach that logically progresses from simplification to practical engagement and collaborative support. Each section builds on the previous one, creating a cohesive argument. Answer1(LightRAG) …
  • Relevance: “Winner”: "Answer2"
    “Explanation”: Both answers are relevant to the question, but Answer2(PankRAG) maintains a sharper focus on practical strategies that directly address how to make theoretical principles accessible. Answer1(LightRAG) …
  • Coherence: “Winner”: "Answer2"
    “Explanation”:Answer2(PankRAG) exhibits a high degree of coherence, with each section clearly connected to the main theme and to each other. The flow of ideas is smooth and logical. Answer1(LightRAG) …
  • Empowerment: “Winner”: "Answer2"
    “Explanation”: Answer2(PankRAG) empowers the reader by providing actionable strategies and emphasizing the importance of mentorship and collaborative learning. It encourages new developers to engage actively with the material. Answer1(LightRAG) …
SubQ 1-1: What are the key theoretical principles from programming language theory? SubQ 2-1: What are effective methods or resources for teaching programming language theory to beginners? SubQ 1-2: What challenges do new developers face in understanding these principles? SubQ 0-1: How can the key theoretical principles be simplified or presented in a way that is more approachable for new developers? SubQ 0-2: What additional support or mentoring could be beneficial for new developers learning these concepts? Query: How can theoretical principles from programming language theory be made accessible to new developers?
Figure 5: A hierarchical resolution path with global awareness of ACQ

### v2

@@ 3.1 Experimental Setup
Table 2: Overall Performance on the SCQ Datasets
@@
GraphRAG 69.56% 86.80% 62.06% 61.18% 68.94% 69.73% 61.57% 60.40% 69.25% 78.27% 61.82% 60.70%
LightRAG 72.40% 86.78% 65.18% 67.00% 67.92% 68.68% 60.99% 58.42% 68.66% 77.73% 63.08% 62.71%
HippoRAG 72.81% 89.84% 73.69% 72.35% 69.18% 75.68% 62.46% 62.19% 70.98% 82.76% 68.01% 67.27%
PankRAG 74.79% 96.53% 87.89% 84.00% 70.62% 88.75% 64.99% 65.35% 72.71% 92.64% 76.44% 74.68%
  • Datasets We follow the standardized preprocessing pipelines from GraphRAG[3] and LightRAG[8] and evaluate our method on six real‑world datasets, encompassing both SCQ and ACQ. We use the Musique [23] and MultiHop-RAG [22] datasets for SCQ. For ACQ, We evaluate our method on the UltraDomin dataset [19], selecting the four domains used (Agriculture, Legal, Computer Science, and Mix)
  • Baselines We compare PankRAG with the three state-of-the-art methods, including NaiveRAG[20], GraphRAG[3], LightRAG[8] and HippoRAG[13] across all datasets.
  • Evaluation Metrics On the SCQ datasets, we employ four metrics from the Ragas framework [4] to assess our method and the baselines: Answer Relevance, Faithfulness, Context Recall, and Context Precision. Lacking ground-truth for the ACQ dataset, we employ GPT-4o-mini as an automated judge. We conduct blinded, pairwise comparisons between Our evaluation involves calculating win rates for our method and baselines across six dimensions. To mitigate positional bias, the presentation order of answers is alternated, and we report the averaged results. We employed GPT-4o-mini for all core tasks. For text embeddings, we utilized the BAAI/bge-m3 model.
Fig. 3: Impact of a pair of hyperparameters in PankRAG on the Mix(left) and Agriculture(right) dataset
@@ 3.2 Overall Performance on the ACQ Datasets
As demonstrated in Table 1, PankRAG consistently surpasses baseline methods across all evaluation metrics and datasets.
PankRAG demonstrates comprehensive superiority, achieving high average win rates across diverse datasets, ranging from 64% to 74%. This dominance extends across all evaluation dimensions, excelling particularly in Relevance (69.8%) and Empowerment (69.7%). This consistent outperformance is a direct result of its core mechanisms—globally-aware pathway resolution and dependency-aware reranking—validating its effectiveness on ACQ.
Fig. 4: Impact of a pair of hyperparameters in PankRAG on the Multihop-RAG dataset[22]
Table 3: Ablation study of PankRAG on ACQ datasets.
Multihop-RAG MuSiQue Average
PankRAG-Pank 67.96% 88.21% 62.94 62.21 67.45% 75.16% 59.90% 58.73% 67.70% 81.69% 61.42% 60.47%
PankRAG 74.79% 96.53% 87.89% 84.00% 70.62% 88.75% 64.99% 65.35% 72.71% 92.64% 76.44% 74.68%
@@ 3.3 Overall Performance on the SCQ Datasets
As demonstrated in Table 2, PankRAG consistently achieves the best performance across all experimental settings. PankRAG establishes its superiority by outperforming baselines by 11-17% on average. This advantage is driven by substantial gains across key metrics, including Faithfulness (10-26%), Context Precision(7-18%), Context Recall (8-19%), and Answer Relevance(2-6%), underscoring its enhanced efficacy in resolving SCQs.
Notably, GraphRAG[3] and LightRAG[8] show minimal improvements in Context Recall and Context Precision (4%-6% over NaiveRAG), indicating negligible enhancement in retrieved context quality. This aligns with the identified challenge that LLMs struggle to interpret hidden relationships or information in SCQ, leading to misaligned retrieval targets. In contrast, PankRAG’s significant improvements in Context Recall and Context Precision suggest effective mitigation of this issue, attributed to its globally-aware pathway resolution generation module and dependency-aware reranking mechanism. It enables LLMs to better capture hidden relationships and thereby improve context relevance. Concurrent enhancements in Faithfulness and Answer Relevance further indicate reduced hallucination risks and higher answer quality, underscoring PankRAG’s comprehensive superiority over existing graph-based RAG methods.
Table 4: Ablation study of PankRAG on ACQ datasets.
@@
w/o Pank w/o Ank w/o Pank w/o Ank w/o Pank w/o Ank w/o Pank w/o Ank
@@
w/o Ank PankRAG w/o Ank PankRAG w/o Ank PankRAG w/o Ank PankRAG
@@ 3.4 Hyperparameter Analysis
The parameters $\alpha$ and $\beta$ (constrained by $\alpha$ + $\beta$ = 1 to ensure the normalized score range) balance the original quality score against a novel similarity metric derived from resolved dependencies. As shown in Figure 3 and Figure 4, Experimental results demonstrate that the optimal parameters for SCQ are $\alpha$ >0.6 $\beta$ <0.4, whereas for ACQ, $\alpha$ = 0.75 and $\beta$ = 0.25 , with both configurations maximizing retrieval performance.
For SCQ, the optimal performance is achieved with $\alpha$ = 0.6 and $\beta$ = 0.4. This configuration represents a critical trade-off. A higher weight on $\alpha$ ($\alpha$ >0.6) results in insufficient reinforcement from sub-question answers, while a lower weight ($\alpha$ <0.6) leads to an over-emphasis that introduces negative hallucinations, degrading performance. The 0.6:0.4 ratio, therefore, strikes the ideal balance between leveraging sub-questions for context validation and mitigating the risk of performance degradation, yielding the best overall retrieval results. For ACQ, our experiments identified a different optimal configuration of $\alpha$ = 0.75 and $\beta$ = 0.25. This setting was determined as it consistently maximized retrieval performance and achieved the highest win rates against baseline models.
@@ 3.5 Ablation Studies
This section’s ablation studies validate the hierarchical contributions of our two core components: globally-aware pathway resolution (Path) and dependency-aware reranking (Ank).
As shown in Table 3 and Table 4, the results demonstrate that the Path module is a foundational component crucial for enhancing retrieval quality. The model variant equipped with Path (PankRAG-Ank) achieves significant improvements in Context Precision and Context Recall over the variant without it (PankRAG-Pank). This finding underscores the necessity of the globally-aware pathway resolution for establishing a high-quality context. Furthermore, the Ank mechanism provides a critical, synergistic enhancement on top of the Path framework. The full PankRAG system, which incorporates the Ank module, shows consistent superiority over the PankRAG-Ank variant across all evaluated metrics. This highlights the critical role of the dependency-aware reranking mechanism in optimizing and refining the final results.
@@ 4 Conclusion
In this paper, we propose PankRAG, a novel RAG framework designed to enhance LLMs’ capacity to accurately interpret and retrieve information for complex queries. PankRAG combines a globally aware pathway resolution module with a dependency‐aware reranking mechanism. By executing a globally coherent reasoning pathway and augmenting retrieved content through dependency‐aware reranking, our framework captures and resolves hidden relationships and information in both SCQ and ACQ, significantly enhancing retrieval performance and response quality.